OpenStudy (anonymous):

Choose the correct answers. Sue Jones is insured for bodily injury in the amount of 10/20. She is at fault in an accident in which Albert Smith is injured. Also injured in the accident is a pedestrian, Sam Dixon. Albert's medical expenses total $12,000, Sam's come to $8,000. How much will the insurance pay? a. for Albert b. for Sam c. total

OpenStudy (ltrout):

Well Sue is insured for 10/20, or 1/2 of the total cost. Albert's + Sam's total cost = 20,000. 1/2 of 20,000 = 10,000. so she can cover for Sam. :)

OpenStudy (anonymous):

thanks as far as albert would it still be 12000?

OpenStudy (ltrout):

I believe I understood that right. It's worded weird.

OpenStudy (ltrout):

Well she can cover for 10,000 total.

OpenStudy (foreveralone01):

He's right. Ugh I always get beaten to answering lol

OpenStudy (anonymous):

there was also a chart but it wouldt let me post it here

OpenStudy (ltrout):

What does the chart show?

OpenStudy (ltrout):

If you've downloaded the chart, just click Attach File under the reply box. If it's online, just right click and click Copy URL address and post that.

OpenStudy (anonymous):

it showed different premium liability coverage.the one that sam fell under was the 10/20 one but im confused at to whether albert would be covered in full from sams nsurance or sam would pay up

OpenStudy (ltrout):

This question is kind of confusing. And um thanks @hotperson lol

OpenStudy (ltrout):

Is this an online schooling or is it a paper?

OpenStudy (tkhunny):

You'll have to explain what "10/20" coverage is.

OpenStudy (anonymous):

there it is lol

OpenStudy (ltrout):

This question is confusing. If I understand right, they are liable for 10/20 of the final cost of the accident?

OpenStudy (anonymous):

yea thats correct but like the question says i need the coverage for sam too that ther insurance will pay

OpenStudy (ltrout):

Wait I get it now. I think. A B C aren't actually choices, they are what you have to find out correct? If so just take 1/2 of each of their prices. 1/2 of albert's 12,000 = 6,000 + 1/2 of sam's 8,000 = 4,000. So a = 6,000, b = 4,000 c = 10,000. I think this is correct now

OpenStudy (anonymous):

yes you got it!! understood that just now as its 1/2 including both thank youu bro

OpenStudy (ltrout):

Haha no problem. Good luck :) Care to medal?

OpenStudy (anonymous):

sure lol but kinda new how so u do that

OpenStudy (ltrout):

You see the blue "Best Response" button next to my answer? Just click that :)

OpenStudy (anonymous):

you lvld up bro lol